What is the total expense for manufacturing 400 tables at Woodlike Inc. if the fixed costs are $1,702,300 and each tables variable cost is $21?


1
Expert's answer
2021-09-24T11:08:42-0400

TC= FC_n + VC_n, n=400, VC=21, FV_n=1702300

VC_n=VC*N=21*400=8400


TC=1702300 + 8400= 1,710,700$